# Hic Sunt Dracones: Uncovering Dynamical Perturbers Within the Habitable Zone

### v2

@@ 1 Introduction
In this paper, we present the results of an analysis of current planetary parameter uncertainties, their effect on orbital dynamics, and the calculated RV signature of potential dynamical perturbers that may yet be lurking within the HZ of the proposed HWO target stars. To borrow from an old Latin phrase, “Hic sunt dracones”, meaning “Here be dragons”, is an apt phrase when exploring the HZ for objects whose presence may render the region unsuitable for habitable environments. In Section 2, we provide an assessment of the derived uncertainties for planet mass and semi-major axis, both of which are key parameters for dynamical simulations, and their dependence on stellar mass and orbital period measurements. In Section 3, we describe a proposed HWO target list and compare the expected RV semi-amplitudes for a Neptune-mass planet in the center of each star’s HZ, along with the challenges in achieving the necessary RV precision to detect such signatures. Section 4 provides an additional discussion regarding the implications of our results and recommendations for a continued dynamical assessment of stars that are considered as future direct imaging targets. A summary of our work and concluding remarks are delivered in Section 5.
2 Derived Planetary System Uncertainties
@@
To investigate the uncertainties in planetary system properties, we extracted data from the Confirmed Planets Table111doi:10.5281/zenodo.11020 of the NASA Exoplanet Archive (Akeson et al., 2013) as accessed on 2024 August 22 (NASA Exoplanet Archive, 2024). We selected the default parameter set for each planet included in the table, and selected only those planets with valid (non-null) values and uncertainties for the semi-major axis ($a$), planet mass ($M_{p}$ or $M_{p}\sin i$), orbital eccentricity ($e$), orbital period ($P$), and the mass of the host star ($M_{\star}$). This provided 1062 planets that passed our criteria for inclusion in our sample. We note that a small number of the planets in this list have masses derived from transit timing variations ($<4$%) rather than radial velocity data but we restrict the overall considerations of this paper to RV-focused efforts as the transit probability for planets approaching the HZ of sun-like stars falls off quickly (Kane & von Braun, 2008).
2.1 Correlations of Uncertainties
@@
Figure 1: Percentage uncertainties for the planetary masses (left column) and semi-major axes (right column) of our confirmed planet list as a function of their respective parameters (top row; panels A and B), stellar mass uncertainties (middle row; panels C and D), and orbital period uncertainties (bottom row; panels E and F). Planets detected only in RV surveys are denoted as cyan circles, while those detected in both RV and transit surveys are denoted as maroon diamonds. The power law relations between stellar and planetary masses and the semi-major axis, orbital period, and stellar mass (see Equations 2 and 4) are evident as diagonal crowdings of points that set the lower limit in the planet mass and semi-major axis uncertainties. For stars with mass uncertainties above $\sim$20%, the stellar mass uncertainty dominates the planet mass determination, rather than the precision of the RV semi-amplitude measurement. Orbital period uncertainties, however, play only a minor role in the planet mass uncertainties. When calculating the semi-major axis, orbital period uncertainties beyond $\sim$10become the dominant source of error but the stellar mass uncertainty does not produce a distinct boundary identifying when it begins dominating the semi-major axis determination.
@@
The semi-major axes and associated uncertainties are shown in panel B of Figure 1. As for the planet mass uncertainties, the semi-major axis uncertainties span several orders of magnitude, falling in the range 0.005–60%, with mean and median values of 4.2% and 2.2%, respectively. For the planets detected in RV surveys only (cyan circles), the data in panels B, E, and F, show a clear capability of the RV surveys to detect longer period planets relative to the transit detections as expected given the decrease in transit probability with increasing orbital period (Kane & von Braun, 2008). The semi-major axis uncertainties show a distinct rise at separations higher than $\sim$5 AU, where the uncertainties become dominated by the those uncertainties associated with the orbital period rather than stellar mass. This correlation is validated by the data distributions shown in panels D and F, where indeed it can be seen that the uncertainties on semi-major axis correlate broadly with those of stellar mass, but correlates strongly with those of orbital period in the high uncertainty regime. The panels of Figure 1 thus provide a broad overview of the relative contributions toward the planet mass and semi-major axis uncertainty budgets, and highlight the importance of improving characterization of host stars and their mass uncertainties. We further provide the range of percentage uncertainty values for each parameter discussed here, along with the mean, median, standard deviation, median absolute deviation (MAD), and interquartile range (IQR) values, in Table 1. The values shown in the the table are for all planet masses, and also for the planet mass groups on either side of 0.1 $M_{J}$ to emphasize the effects of planet mass on the measurement precision. Note that the high uncertainty for semi-major axis and orbital period for the high mass group ($>0.1$ $M_{J}$) is the result of the observational bias for high mass planets at large separations.
Table 1: Parameter percentage uncertainty statistics.
@@
footnotetext: $P$ uncertainties of zero correspond to uncertainties $<10^{-4}$ days.
@@
2.2 Effects on Dynamical Simulations
As described in Section 2.1, an impetus for improving the planet mass and semi-major axis values lies in their critical role in assessing the dynamical stability and evolution of planetary systems (Maire et al., 2023). Orbital dynamics has proven to be a useful tool in predicting the possible locations of additional planets in known systems (Barnes & Raymond, 2004; Rivera & Haghighipour, 2007; Kopparapu et al., 2009; Agnew et al., 2019), as well as assessing their dynamical packing (Fang & Margot, 2013; Obertas et al., 2017; Kane et al., 2020; Kane, 2023a; Obertas et al., 2023). Dynamical simulations have also been applied to evaluating dynamical stability in the HZ, with application to terrestrial planet climates and the sustainability of temperate conditions (Williams & Pollard, 2002; Dressing et al., 2010; Kane & Gelino, 2012b; Carrera et al., 2016; Kane & Blunt, 2019; Kane et al., 2022; Li et al., 2022; Vervoort et al., 2022). Such work relies upon the robustness of reported planetary system parameters, both in terms of the measurement precision and accounting for other possible planets and interactions (Beaugé et al., 2008; Wittenmyer et al., 2019).
The precise effects of system uncertainties on dynamical analyses are highly dependent upon the system architecture under consideration (Hadden & Lithwick, 2018; Kane, 2023b). By way of example, consider the HD 141399 system, which contains four giant planets with a diversity of star–planet separations (Vogt et al., 2014). The dynamical analysis of the HD 141399 system provided by Kane (2023b) revealed the complex mean motion resonance (MMR) structure of the system, and the consequences for dynamical stability within the HZ. Many of the adjacent primary MMR locations are separated by a distance of only 10%, such as the 3:1 (0.201 AU) and 5:2 (0.227 AU) MMR locations for planet b. As shown in Section 2.1, while the median semi-major axis uncertainty is only 2.2%, values can be as high as 60%. These semi-major axis uncertainties can have significant consequences for the correct diagnosis of dynamical aspects, such as the apsidal behavior of orbital evolution (libration or circulation) near MMR locations Barnes & Greenberg (2006); Kane & Raymond (2014) and can ultimately lead to incorrect conclusions regarding the long-term stability prognosis of planetary systems.
3 Perturbers Within the Habitable Zone
@@
3.1 Habitable Worlds Observatory Targets
@@
The recent work by Kane et al. (2024) demonstrated the importance of understanding the full dynamical effects of system architectures on the possible presence of planets in the HZ. Here, we use the proposed HWO target list provided by Mamajek & Stapelfeldt (2024) to calculate the signature of an additional, Neptune-mass, planet within the HZ (described in Section 3.2) for all 164 target stars. The stars are all relatively bright, with a V magnitude range of 0.0–7.4, and span a broad range of spectral types, with an effective temperature range of 3601–6990 K.
@@
3.2 Required Radial Velocity Observations
There are a variety of planet mass and orbits that can cause significant impact orbital stability within the HZ, from those whose orbits lie entirely within the HZ to giant planets on much longer period but more eccentric orbits (e.g., Kane & Blunt, 2019). Here, we consider the simple case of a Neptune-mass planet ($M_{p}=17$ $M_{\oplus}$) on a circular orbit in the middle of the HZ. This particular configuration was chosen based on the recent re-analysis of the GJ 357 system by Kane & Fetherolf (2023), which showed that even a 10 $M_{\oplus}$ mass planet in the HZ can render a substantial fraction of that region unstable for other terrestrial planets. For each of the 164 proposed HWO targets stars (see Section 3.1), we calculated the OHZ boundaries and the RV semi-amplitude for a Neptune-mass planet that lies halfway between them. The results of these calculations are shown in Figure 3, both as a function of the host star effective temperature (left panel) and as a histogram (right panel). As expected, the RV semi-amplitude decreases with increasing stellar effective temperature, since both the stellar mass and semi-major axis of the HZ boundaries increase in that direction. There are two M dwarfs (GJ 411 and GJ 887) for which the RV semi-amplitude exceeds 6 m s-1, but the vast majority have a RV semi-amplitude that is less than 5 m s-1.
@@ 3.2 Required Radial Velocity Observations
Recovering long period, small RV semi-amplitude signatures remains challenging in the face of stellar activity and correlated noise (Narayan et al., 2005; Kane et al., 2009; Wittenmyer et al., 2013; Fischer et al., 2016; Langellier et al., 2021; Luhn et al., 2023). For example, the community-wide RV challenge conducted by Dumusque et al. (2017) used data acquired with the High Accuracy Radial velocity Planet Searcher (HARPS) spectrograph (Pepe et al., 2000), into which planetary signatures were injected. The results of that challenge indicate that, for a star with 1.5 m s-1stellar jitter, a planetary signature with a RV amplitude of 1.5 m s-1 would require $\sim$60 HARPS observations to achieve a robust detection, depending on the brightness of the host star, the observing cadence adopted, and the precision of the specific instrument used. Indeed, recent studies of archival RV measurements by Laliotis et al. (2023) and Harada et al. (2024b) demonstrated that the RV sensitivity is not yet sufficient to extract the signature of Saturn-mass planets that may be present in the HZ for many of the proposed HWO target stars. Additional effort will be required for a more detailed assessment of chromospheric activity (Cale et al., 2021; Isaacson et al., 2024), particularly for the coolest host stars where the HZ orbits and overlap with stellar rotation periods (Vanderburg et al., 2016). Although a full diagnosis of the specific number of observations and optimum scheduling, reduction, and analysis strategies required for all HWO stars is a complex and multi-dimensional problem, it is clear that significant observational resources are required to adequately uncover the signatures of potential gravitational perturbers that yet remain in or near the HZ of those systems. We provide specific recommendations toward achieving the goal of detecting such planets in Section 5.
4 Discussion
@@
There are several additional points of consideration to add to the details of this paper. Since the discussion here has been largely focused on the planetary signatures and uncertainties that arise from RV observations, there remains the problem of an unknown orbital inclination in the majority of these cases. The RV semi-amplitude calculations of Section 3.2 assume an edge-on inclination, which will not be true in the vast majority of cases, resulting in more stringent precision requirements to detect the correspondingly smaller amplitude planetary RV signatures. Orbital dynamics for multi-planet systems may be used to place lower limits on the inclination of the planetary orbits, and subsequent effects on the HZ (Kane & Gelino, 2014; Korolik et al., 2023). The coplanarity of the orbits within a system is difficult to determine for non-transiting systems (Fang & Margot, 2012; Ballard & Johnson, 2016), however, and will also influence the calculated dynamics of the HZ (Kane, 2016). A further pathway toward resolving the inclination ambiguity is the use of astrometric observations, in particular those of the Gaia mission (Perryman et al., 2014; Gaia Collaboration et al., 2021; Feng et al., 2021) in combination with RV data (Brandt et al., 2019; Kiefer et al., 2021; Winn, 2022). Such methodology may only be applied for those stars that are fainter than the Gaia bright limit of $G\approx 3$ (Lindegren et al., 2018). Though Gaia is unlikely to reveal planets within the HZ, the constraints placed on the inclination of outer, giant planetary companions can provide valuable insight into the inclination and dynamics of inner planets within the system. The use of Gaia data has the additional benefit of greatly improving stellar mass uncertainties (Stassun et al., 2017; Fulton et al., 2017), and thus increasing the reliability of dynamical simulations, as discussed in Section 2. However, the inevitable outcome of such observations will be to reveal that the true planetary masses are indeed larger than the minimum masses inferred from the RV data alone, further inhibiting the dynamical real estate available with the HZ of target systems.
@@ 5 Conclusions
First, we recommend that the exoplanet community invest in an update of the stellar parameters for all known exoplanet host stars using a self-consistent and homogeneous analysis. One potential approach is to fit the stars’ spectral energy distributions, combined with precise distance estimates from Gaia DR3, and then using Bayesian Model Averaging to incorporate information from multiple stellar models (see, e.g., Vines & Jenkins, 2022; Harada et al., 2024c). Alternatively, high resolution spectra for these stars can be used to derive spectroscopic stellar parameters, either via comparisons to a grid of spectral templates (e.g., Petigura et al., 2017) or via direct spectral synthesis that computes radiative transfer through a model stellar atmosphere (e.g., Valenti & Piskunov, 1996; Soto & Jenkins, 2018). Where possible, the spectroscopically derived parameters should be used as constraints to interpolate upon stellar isochrones (e.g., Demarque et al., 2004; Choi et al., 2016) to derive additional mass, radius, and age ranges for the stars. As most exoplanet characteristics are derived parameters whose value and uncertainty rely on our knowledge of the host star’s mass and/or radius, updates to the stellar parameters will then need to be propagated through to the planet parameters. A full re-analysis of the original time series data that led to the detection of the planet, either the transit photometry or RV time series, would provide the most cohesive view of each system. However, an intermediate option would be to take the measured quantity (either the transit depth or the RV semi-amplitude) and re-derive the corresponding exoplanet radius or mass measurement using the updated stellar radius or mass.
@@
Second, we recommend the initiation of dedicated RV surveys that target the list of the likely HWO direct imaging targets and search for evidence of disruptive planets. Recent simulation efforts from Luhn et al. (2023) and observational efforts of the sun from Klein et al. (2024) and Ford et al. (2024) suggest that the detection of HZ Neptunes and even super-Earths at the $\sim$10 Earth mass level is attainable with modern RV instruments. The application of a Fischer information metric that incorporates information about correlated stellar noise (see, e.g., Cloutier et al., 2018; Gupta & Bedell, 2024; Lam et al., 2024) may provide an efficient way to survey the HZs of these stars. The design and execution of such a survey will likely require multiple teams working in coordination, as the HWO targets span the entire sky and cannot be comprehensively observed from a single facility. Such a survey will also require a multi-year investment of observing resources as the orbital periods of these HWO HZs are 100s of days long, and establishing the stability of an RV signal over multiple orbits is often a requirement for confirming its Keplerian nature, particularly for those RV semi-amplitudes near the 1.5 m s-1peak described in Section 3.2. As a first step, we therefore recommend that the exoplanet community determine the most effective and efficient way to conduct such a survey and compare the resulting observational needs with the current suite of accessible RV facilities to determine what additional resources / observing support may be required. Once such a plan has been developed and vetted by the community, execution should begin as soon as possible to ensure that the outer regions of the HZ are thoroughly investigated in time to provide insights to HWO’s architectural design and science operations. The detection of these dynamical perturbers within the HZ will not only guide the target selection of direct imaging missions but may also guide the mission architectures if such perturbers are found to be common place, thus requiring an increase in mission target lists and a corresponding increase in mission aperture or duration. The results of such a survey would also provide fundamental demographics information that informs the occurrence of potentially habitable planets.

### v3

@@ I Introduction
In this paper, we present the results of an analysis of current planetary parameter uncertainties, their effect on orbital dynamics, and the calculated RV signature of potential dynamical perturbers that may yet be lurking within the HZ of the proposed HWO target stars. To borrow from an old Latin phrase, “Hic sunt dracones”, meaning “Here be dragons”, is an apt phrase when exploring the HZ for objects whose presence may render the region unsuitable for habitable environments. In Section II, we provide an assessment of the derived uncertainties for planet mass and semi-major axis, both of which are key parameters for dynamical simulations, and their dependence on stellar mass and orbital period measurements. In Section III, we describe a proposed HWO target list and compare the expected RV semi-amplitudes for a Neptune-mass planet in the center of each star’s HZ, along with the challenges in achieving the necessary RV precision to detect such signatures. Section IV provides an additional discussion regarding the implications of our results and recommendations for a continued dynamical assessment of stars that are considered as future direct imaging targets. A summary of our work and concluding remarks are delivered in Section V.
II Derived Planetary System Uncertainties
@@
To investigate the uncertainties in planetary system properties, we extracted data from the Confirmed Planets Table111doi:10.5281/zenodo.11020 of the NASA Exoplanet Archive (Akeson et al., 2013) as accessed on 2024 August 22 (NASA Exoplanet Archive, 2024). We selected the default parameter set for each planet included in the table, and selected only those planets with valid (non-null) values and uncertainties for the semi-major axis ($a$), planet mass ($M_{p}$ or $M_{p}\sin i$), orbital eccentricity ($e$), orbital period ($P$), and the mass of the host star ($M_{\star}$). This provided 1062 planets that passed our criteria for inclusion in our sample. We note that a small number of the planets in this list have masses derived from transit timing variations (<4<4%) rather than radial velocity data but we restrict the overall considerations of this paper to RV-focused efforts as the transit probability for planets approaching the HZ of sun-like stars falls off quickly (Kane & von Braun, 2008).
II.1 Correlations of Uncertainties
@@
Figure 1: Percentage uncertainties for the planetary masses (left column) and semi-major axes (right column) of our confirmed planet list as a function of their respective parameters (top row; panels A and B), stellar mass uncertainties (middle row; panels C and D), and orbital period uncertainties (bottom row; panels E and F). Planets detected only in RV surveys are denoted as cyan circles, while those detected in both RV and transit surveys are denoted as maroon diamonds. The power law relations between stellar and planetary masses and the semi-major axis, orbital period, and stellar mass (see Equations 2 and 4) are evident as diagonal crowdings of points that set the lower limit in the planet mass and semi-major axis uncertainties. For stars with mass uncertainties above $\sim$20%, the stellar mass uncertainty dominates the planet mass determination, rather than the precision of the RV semi-amplitude measurement. Orbital period uncertainties, however, play only a minor role in the planet mass uncertainties. When calculating the semi-major axis, orbital period uncertainties beyond $\sim$10% become the dominant source of error but the stellar mass uncertainty does not produce a distinct boundary identifying when it begins dominating the semi-major axis determination.
@@
The semi-major axes and associated uncertainties are shown in panel B of Figure 1. As for the planet mass uncertainties, the semi-major axis uncertainties span several orders of magnitude, falling in the range 0.005–60%, with mean and median values of 4.2% and 2.2%, respectively. For the planets detected in RV surveys only (cyan circles), the data in panels B, E, and F, show a clear capability of the RV surveys to detect longer period planets relative to the transit detections as expected given the decrease in transit probability with increasing orbital period (Kane & von Braun, 2008). The semi-major axis uncertainties show a distinct rise at separations higher than $\sim$5 AU, where the uncertainties become dominated by the those uncertainties associated with the orbital period rather than stellar mass. This correlation is validated by the data distributions shown in panels D and F, where indeed it can be seen that the uncertainties on semi-major axis correlate broadly with those of stellar mass, but correlates strongly with those of orbital period in the high uncertainty regime. The panels of Figure 1 thus provide a broad overview of the relative contributions toward the planet mass and semi-major axis uncertainty budgets, and highlight the importance of improving characterization of host stars and their mass uncertainties. We further provide the range of percentage uncertainty values for each parameter discussed here, along with the mean, median, standard deviation, median absolute deviation (MAD), and interquartile range (IQR) values, in Table 1. The values shown in the the table are for all planet masses, and also for the planet mass groups on either side of 0.1 $M_{J}$ to emphasize the effects of planet mass on the measurement precision. Note that the high uncertainty for semi-major axis and orbital period for the high mass group (>0.1>0.1 $M_{J}$) is the result of the observational bias for high mass planets at large separations.
Table 1: Parameter percentage uncertainty statistics.footnotetext: $P$ uncertainties of zero correspond to uncertainties $10^{-4}$ days.
@@
II.2 Effects on Dynamical Simulations
As described in Section II.1, an impetus for improving the planet mass and semi-major axis values lies in their critical role in assessing the dynamical stability and evolution of planetary systems (Maire et al., 2023). Orbital dynamics has proven to be a useful tool in predicting the possible locations of additional planets in known systems (Barnes & Raymond, 2004; Rivera & Haghighipour, 2007; Kopparapu et al., 2009; Agnew et al., 2019), as well as assessing their dynamical packing (Fang & Margot, 2013; Obertas et al., 2017; Kane et al., 2020; Kane, 2023a; Obertas et al., 2023). Dynamical simulations have also been applied to evaluating dynamical stability in the HZ, with application to terrestrial planet climates and the sustainability of temperate conditions (Williams & Pollard, 2002; Dressing et al., 2010; Kane & Gelino, 2012b; Carrera et al., 2016; Kane & Blunt, 2019; Kane et al., 2022; Li et al., 2022; Vervoort et al., 2022). Such work relies upon the robustness of reported planetary system parameters, both in terms of the measurement precision and accounting for other possible planets and interactions (Beaugé et al., 2008; Wittenmyer et al., 2019).
The precise effects of system uncertainties on dynamical analyses are highly dependent upon the system architecture under consideration (Hadden & Lithwick, 2018; Kane, 2023b). By way of example, consider the HD 141399 system, which contains four giant planets with a diversity of star–planet separations (Vogt et al., 2014). The dynamical analysis of the HD 141399 system provided by Kane (2023b) revealed the complex mean motion resonance (MMR) structure of the system, and the consequences for dynamical stability within the HZ. Many of the adjacent primary MMR locations are separated by a distance of only 10%, such as the 3:1 (0.201 AU) and 5:2 (0.227 AU) MMR locations for planet b. As shown in Section II.1, while the median semi-major axis uncertainty is only 2.2%, values can be as high as 60%. These semi-major axis uncertainties can have significant consequences for the correct diagnosis of dynamical aspects, such as the apsidal behavior of orbital evolution (libration or circulation) near MMR locations Barnes & Greenberg (2006); Kane & Raymond (2014) and can ultimately lead to incorrect conclusions regarding the long-term stability prognosis of planetary systems.
III Perturbers Within the Habitable Zone
@@
III.1 Habitable Worlds Observatory Targets
@@
The recent work by Kane et al. (2024) demonstrated the importance of understanding the full dynamical effects of system architectures on the possible presence of planets in the HZ. Here, we use the proposed HWO target list provided by Mamajek & Stapelfeldt (2024) to calculate the signature of an additional, Neptune-mass, planet within the HZ (described in Section III.2) for all 164 target stars. The stars are all relatively bright, with a V magnitude range of 0.0–7.4, and span a broad range of spectral types, with an effective temperature range of 3601–6990 K.
@@
III.2 Required Radial Velocity Observations
There are a variety of planet mass and orbits that can cause significant impact orbital stability within the HZ, from those whose orbits lie entirely within the HZ to giant planets on much longer period but more eccentric orbits (e.g., Kane & Blunt, 2019). Here, we consider the simple case of a Neptune-mass planet ($_{p}=17 MM_{\oplus}$) on a circular orbit in the middle of the HZ. This particular configuration was chosen based on the recent re-analysis of the GJ 357 system by Kane & Fetherolf (2023), which showed that even a 10 $M_{\oplus}$ mass planet in the HZ can render a substantial fraction of that region unstable for other terrestrial planets. For each of the 164 proposed HWO targets stars (see Section III.1), we calculated the OHZ boundaries and the RV semi-amplitude for a Neptune-mass planet that lies halfway between them. The results of these calculations are shown in Figure 3, both as a function of the host star effective temperature (left panel) and as a histogram (right panel). As expected, the RV semi-amplitude decreases with increasing stellar effective temperature, since both the stellar mass and semi-major axis of the HZ boundaries increase in that direction. There are two M dwarfs (GJ 411 and GJ 887) for which the RV semi-amplitude exceeds 6 m s-1, but the vast majority have a RV semi-amplitude that is less than 5 m s-1.
@@ III.2 Required Radial Velocity Observations
Recovering long period, small RV semi-amplitude signatures remains challenging in the face of stellar activity and correlated noise (Narayan et al., 2005; Kane et al., 2009; Wittenmyer et al., 2013; Fischer et al., 2016; Langellier et al., 2021; Luhn et al., 2023). For example, the community-wide RV challenge conducted by Dumusque et al. (2017) used data acquired with the High Accuracy Radial velocity Planet Searcher (HARPS) spectrograph (Pepe et al., 2000), into which planetary signatures were injected. The results of that challenge indicate that, for a star with 1.5 m s-1stellar jitter, a planetary signature with a RV amplitude of 1.5 m s-1 would require $\sim$60 HARPS observations to achieve a robust detection, depending on the brightness of the host star, the observing cadence adopted, and the precision of the specific instrument used. Indeed, recent studies of archival RV measurements by Laliotis et al. (2023) and Harada et al. (2025) demonstrated that the RV sensitivity is not yet sufficient to extract the signature of Saturn-mass planets that may be present in the HZ for many of the proposed HWO target stars. Additional effort will be required for a more detailed assessment of chromospheric activity (Cale et al., 2021; Isaacson et al., 2024), particularly for the coolest host stars where the HZ orbits and overlap with stellar rotation periods (Vanderburg et al., 2016). Although a full diagnosis of the specific number of observations and optimum scheduling, reduction, and analysis strategies required for all HWO stars is a complex and multi-dimensional problem, it is clear that significant observational resources are required to adequately uncover the signatures of potential gravitational perturbers that yet remain in or near the HZ of those systems. We provide specific recommendations toward achieving the goal of detecting such planets in Section V.
IV Discussion
@@
There are several additional points of consideration to add to the details of this paper. Since the discussion here has been largely focused on the planetary signatures and uncertainties that arise from RV observations, there remains the problem of an unknown orbital inclination in the majority of these cases. The RV semi-amplitude calculations of Section III.2 assume an edge-on inclination, which will not be true in the vast majority of cases, resulting in more stringent precision requirements to detect the correspondingly smaller amplitude planetary RV signatures. Orbital dynamics for multi-planet systems may be used to place lower limits on the inclination of the planetary orbits, and subsequent effects on the HZ (Kane & Gelino, 2014; Korolik et al., 2023). The coplanarity of the orbits within a system is difficult to determine for non-transiting systems (Fang & Margot, 2012; Ballard & Johnson, 2016), however, and will also influence the calculated dynamics of the HZ (Kane, 2016). A further pathway toward resolving the inclination ambiguity is the use of astrometric observations, in particular those of the Gaia mission (Perryman et al., 2014; Gaia Collaboration et al., 2021; Feng et al., 2021) in combination with RV data (Brandt et al., 2019; Kiefer et al., 2021; Winn, 2022). Such methodology may only be applied for those stars that are fainter than the Gaia bright limit of $G\approx 3$ (Lindegren et al., 2018). Though Gaia is unlikely to reveal planets within the HZ, the constraints placed on the inclination of outer, giant planetary companions can provide valuable insight into the inclination and dynamics of inner planets within the system. The use of Gaia data has the additional benefit of greatly improving stellar mass uncertainties (Stassun et al., 2017; Fulton et al., 2017), and thus increasing the reliability of dynamical simulations, as discussed in Section II. However, the inevitable outcome of such observations will be to reveal that the true planetary masses are indeed larger than the minimum masses inferred from the RV data alone, further inhibiting the dynamical real estate available with the HZ of target systems.
@@ V Conclusions
First, we recommend that the exoplanet community invest in an update of the stellar parameters for all known exoplanet host stars using a self-consistent and homogeneous analysis. One potential approach is to fit the stars’ spectral energy distributions, combined with precise distance estimates from Gaia DR3, and then using Bayesian Model Averaging to incorporate information from multiple stellar models (see, e.g., Vines & Jenkins, 2022; Harada et al., 2024b). Alternatively, high resolution spectra for these stars can be used to derive spectroscopic stellar parameters, either via comparisons to a grid of spectral templates (e.g., Petigura et al., 2017) or via direct spectral synthesis that computes radiative transfer through a model stellar atmosphere (e.g., Valenti & Piskunov, 1996; Soto & Jenkins, 2018). Where possible, the spectroscopically derived parameters should be used as constraints to interpolate upon stellar isochrones (e.g., Demarque et al., 2004; Choi et al., 2016) to derive additional mass, radius, and age ranges for the stars. As most exoplanet characteristics are derived parameters whose value and uncertainty rely on our knowledge of the host star’s mass and/or radius, updates to the stellar parameters will then need to be propagated through to the planet parameters. A full re-analysis of the original time series data that led to the detection of the planet, either the transit photometry or RV time series, would provide the most cohesive view of each system. However, an intermediate option would be to take the measured quantity (either the transit depth or the RV semi-amplitude) and re-derive the corresponding exoplanet radius or mass measurement using the updated stellar radius or mass.
@@
Second, we recommend the initiation of dedicated RV surveys that target the list of the likely HWO direct imaging targets and search for evidence of disruptive planets. Recent simulation efforts from Luhn et al. (2023) and observational efforts of the sun from Klein et al. (2024) and Ford et al. (2024) suggest that the detection of HZ Neptunes and even super-Earths at the $\sim$10 Earth mass level is attainable with modern RV instruments. The application of a Fischer information metric that incorporates information about correlated stellar noise (see, e.g., Cloutier et al., 2018; Gupta & Bedell, 2024; Lam et al., 2024) may provide an efficient way to survey the HZs of these stars. The design and execution of such a survey will likely require multiple teams working in coordination, as the HWO targets span the entire sky and cannot be comprehensively observed from a single facility. Such a survey will also require a multi-year investment of observing resources as the orbital periods of these HWO HZs are 100s of days long, and establishing the stability of an RV signal over multiple orbits is often a requirement for confirming its Keplerian nature, particularly for those RV semi-amplitudes near the 1.5 m s-1peak described in Section III.2. As a first step, we therefore recommend that the exoplanet community determine the most effective and efficient way to conduct such a survey and compare the resulting observational needs with the current suite of accessible RV facilities to determine what additional resources / observing support may be required. Once such a plan has been developed and vetted by the community, execution should begin as soon as possible to ensure that the outer regions of the HZ are thoroughly investigated in time to provide insights to HWO’s architectural design and science operations. The detection of these dynamical perturbers within the HZ will not only guide the target selection of direct imaging missions but may also guide the mission architectures if such perturbers are found to be common place, thus requiring an increase in mission target lists and a corresponding increase in mission aperture or duration. The results of such a survey would also provide fundamental demographics information that informs the occurrence of potentially habitable planets.